OpenStudy (anonymous):

Determine if triangle ABC with coordinates A (0, 2), B (2, 5), and C (-1, 7) is an isosceles triangle. Use evidence to support your claim. If it is not an isosceles triangle, what changes can be made to make it isosceles? Be specific.

OpenStudy (tkhunny):

Can you calculate the distance between each of the three points?

OpenStudy (anonymous):

Well yeah I can do that, but how can you classify a triangle a Isosceles triangle?

OpenStudy (tkhunny):

Why don't you do it and see if anything jumps out at you. What the definition of an isosceles triangle?

OpenStudy (anonymous):

\[ AB=\sqrt{13}\\ AC=\sqrt{26}\\ BC=\sqrt{13} \] Which two sides are equal? Can you conclude the this triangle is isosceles ?

OpenStudy (tkhunny):

Are there TWO equal sides? You're done. It's isosceles.